OpenStudy (anonymous):

Find the area bounded by the x-axis, y= sqrt(x-2), and y=-x

OpenStudy (anonymous):

first we need to find the limits if integration, we do this be setting the equations equal to each other

OpenStudy (anonymous):

I found that this needs to be integrated by parts (if you integrate horizontally). The first limits are from -2 to 0, and the other limit is from 0 to 2

OpenStudy (anonymous):

okay so simply integrate \[(\sqrt{x-2})-(-x)\], using those limits

OpenStudy (anonymous):

For the first area between -2 and 0, I got \[4^{3/2} / 3\] I was wondering if you get the same. This was a question that we did in class, and I got an answer different from the prof. Also, I used U-substitution and ended up changing the limits of integration. The prof. did the problem a different way that I've never seen before. So really, I'm just wondering if you get the same answer for that first area.

OpenStudy (anonymous):

i am not in a position to do the computation right now, but let me get you someone who can alright

OpenStudy (anonymous):

k

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

there is no solution to this. you have \[y=\sqrt{x-2}\] and \[y=-x\] if i read it correctly. these do not intersect

OpenStudy (anonymous):

They intersect at (-2,2). But, this is also bounded by the x-axis. Because of this, you need to integrate by parts, or integrate with respect to y (which I'm not supposed to do yet)

OpenStudy (anonymous):

no they don't

OpenStudy (anonymous):

oh crud, I meant sqrt(2-x). I'm sorry I didn't catch that

OpenStudy (anonymous):

in particular you cannot evaluate \[f(x)=\sqrt{x-2}\] at -2 you will get \[\sqrt{-4}\]

OpenStudy (anonymous):

oooooooooooooooooh ok maybe we can do it now

OpenStudy (anonymous):

now they intersect at -2

OpenStudy (anonymous):

so you have to break up the integral into two pieces right?

OpenStudy (anonymous):

because you are above the x - axis. so first take \[\int_{-2}^0(\sqrt{2-x}+x)dx\]then \[\int_0^2\sqrt{2-x}dx\]

OpenStudy (anonymous):

first anti derivative is \[-\frac{2}{3}(2-x)^{\frac{3}{2}}+\frac{x^2}{2}\] second is just \[-\frac{2}{3}(2-x)^{\frac{3}{2}}\]

OpenStudy (anonymous):

now plug in limits and be done

OpenStudy (anonymous):

My teacher said we needed to use u-substitution

OpenStudy (anonymous):

fine. you can use \[u=2-x\] \[du=-dx\] \[-du=dx\] and get \[\int\sqrt{2-x}dx=-\int\sqrt{u}du=-\frac{2}{3}u^{\frac{3}{2}}\]

OpenStudy (anonymous):

that is what i did but i didn't make it explicit. only off by a minus sign that is why

OpenStudy (anonymous):

now replace u by 2-x and get the answer i wrote. (i hope)

OpenStudy (anonymous):

alright, that's what I was looking for. Thanks!

OpenStudy (anonymous):

yw